Differentiate what type of return to scale the following production function exhibit and explain your reason

            A. Q=L0.5 K0.5                               C. Q = L0.3 K0.5

            B. Q=L2/3 K2/3                   D. Q=ALa Kb

1
Expert's answer
2021-10-03T14:14:19-0400

A. Q=L0.5 K0.5                            

  

"0.5+0.5=1"

proportionally increases


            B. Q=L2/3 K2/3   


"\frac{2}{3}+\frac{2}{3}=\frac{4}{3}>1"

disproportionally increases


C. Q = L0.3 K0.5

"0.3+0.5=0.8<1"

disproportionally decreases


   D. Q=ALa Kb

Depends on the value of the sum a + b